An object is thrown vertically downward with speed 20m/s from a height 100m. The time of flight(in s) of the object is (g= 10m/s2) ?
 

Grade:12th pass

1 Answers

Mahima Kanawat
5 years ago
Dear student 
Using second equation of motion
S =ut +1/2at2
Taking downwards negative and upwards positive 
-100 = - 20t + 1/2(-10)t2
-100 = - 20t - 5t2
20 = 4t + 5t2
5t2 +4t - 20 = 0
On solving t = 1.63 seconds 
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT